OpenStudy (pokemon23):

factor completely 18m^2-

OpenStudy (pokemon23):

18n^2-8

OpenStudy (pokemon23):

i got 2(9m^2-4)

OpenStudy (lgbasallote):

2(3m+2)(3m-2)

OpenStudy (bahrom7893):

OK

OpenStudy (lgbasallote):

2(9m^2 - 4) has difference of two squares see?

OpenStudy (bahrom7893):

come on lgba.. i wanted him to finish

OpenStudy (anonymous):

\[(3\sqrt{2}m+2\sqrt{2})(3\sqrt{2}m-2\sqrt{2})\]

OpenStudy (pokemon23):

so I forgot to do one step -_-'

OpenStudy (pokemon23):

so if the question was 2x^2-18n=2(x^2-9)

OpenStudy (pokemon23):

the answe will be 2(x-9) (x+2)

OpenStudy (lgbasallote):

no...2(x^2 -9n) you have an n with 18

OpenStudy (lgbasallote):

and that can no longer be simplified

OpenStudy (pokemon23):

OHH

OpenStudy (anonymous):

\[(3\sqrt{2}m+2\sqrt{2})(3\sqrt{2}m-2\sqrt{2})=\sqrt{2}(3m+2)(3m-2)\] why is that different

OpenStudy (anonymous):

ohh I see sqrt2*sqrt2=2